# Magnetism and hidden quantum geometry in charge neutral twisted trilayer graphene

### v1

@@ Methods
Mirror-symmetric TTG is constructed by aligning the top and bottom layers and twisting the middle layer by a given angle $\theta$. If $\theta$ belongs to a set of commensurate twist angles, moiré primitive vectors are $\vec{L}_{1}^{(m)}=m\vec{a}_{1}-(m+1)\vec{a}_{2}$, $\vec{L}_{2}^{(m)}=-(m+1)\vec{a}_{1}-(2m+1)\vec{a}_{2}$, where $m$ is an integer, which is related to the twist angle $\theta$ by $\cos(\theta_{m})=(3m^{2}+3m+1/2)/(3m^{2}+3m+1)$. Here we use $\theta=1.55^{\circ}$ ($m=21$). The number of atoms in moiré unit cell is $N_{\rm at}=6(3m(m+1)+1)$, which here gives 8322 atoms. We account for the lattice relaxation effects along the z-axis by varying the inter-layer distance based on the stacking configuration [56, 20, 40]. We treat the middle layer as rigid and relax only the top and bottom layers. More details of the relaxation are presented in SM.
To describe the electronic properties of TTG, we employ a tight-binding model for $p_{z}$ atomic orbitals defined by Eq. (1) and (2). We use the following parameters $a=1.412$, $c=3.36$, $\gamma_{0}=-2.835$ eV, $\gamma_{1}=0.48$ eV. Here, $a$ is the carbon-carbon distance, $c$ is the inter-layer distance of an unrelaxed system, $\gamma_{0}(\gamma_{1})$ is the value of the nearest in-plane (out-of-plane) neighbor hopping. The dimensionless decay constants are $\lambda_{1}=3.15$ in-plane and $\lambda_{2}=7.50$ out-of-plane. A more detailed description of the model can be found in Refs. [57, 52].
@@
Finally, we calculate the multiband Chern numbers using a numerical procedure on a discretized Brillouin zone [58]. We define $n$-bands link variable $U_{\mu}(\vec{k}_{i})$ from the chosen set of wave functions $\psi=(\ket{\phi_{1}},...,\ket{\phi_{n}})$ as $U_{\mu}(\vec{k}_{i})\equiv\det\psi^{\dagger}(\vec{k}_{i})\psi(\vec{k}_{i}+\hat{\mu})/|\det\psi^{\dagger}(\vec{k}_{i})\psi(\vec{k}_{i}+\hat{\mu})|$. Here, $\mu=1,2$ represents the possible directions in 2D while $\hat{\mu}$ is the vector connecting to the nearest $i$-th $k$ point with its nearest neighbor in direction $\mu$. From the link variable, we can define mBZ as $F(\vec{k}_{i})=\ln\left(U_{1}(\vec{k}_{i})U_{2}(\vec{k}_{i}+\hat{1})U_{1}(\vec{k}_{i}+\hat{2})^{-1}U_{2}(\vec{k}_{i})^{-1}\right)$. The Chern number for the $n$ bands is then defined $c_{\psi}\equiv(2\pi i)^{-1}\sum_{i}F(\vec{k}_{i})$. Note that since $U_{\mu}(\vec{k}_{i})$ is normalized, $F(\vec{k}_{i})$ is purely imaginary, and the Chern number $c_{\psi}$ is real.
@@ Acknowledgments
A.W.R., D.M. and P.H. were supported by NSERC Discovery Grant No. RGPIN 2019-05714, NSERC Alliance Quantum Grant No. ALLRP/578466-2022, the QSP-078 project of the Quantum Sensors Program at the National Research Council of Canada, University of Ottawa Research Chair in Quantum Theory of Materials, Nanostructures, and Devices. M.B. has been supported by the National Science Centre, Poland, under Grant No. 2021/43/D/ST3/01989. This paper was partly enabled by support provided by the Digital Research Alliance of Canada [59].

### v2

@@ Methods
Mirror-symmetric TTG is constructed by aligning the top and bottom layers and twisting the middle layer by a given angle $\theta$. If $\theta$ belongs to a set of commensurate twist angles, moiré primitive vectors are $\vec{L}_{1}^{(m)}=m\vec{a}_{1}-(m+1)\vec{a}_{2}$, $\vec{L}_{2}^{(m)}=-(m+1)\vec{a}_{1}-(2m+1)\vec{a}_{2}$, where $m$ is an integer, which is related to the twist angle $\theta$ by $\cos(\theta_{m})=(3m^{2}+3m+1/2)/(3m^{2}+3m+1)$. Here we use $\theta=1.55^{\circ}$ ($m=21$). The number of atoms in moiré unit cell is $N_{\rm at}=6(3m(m+1)+1)$, which here gives 8322 atoms. We account for the lattice relaxation effects along the z-axis by varying the inter-layer distance based on the stacking configuration [Lopez-Bezanilla_Lado_2020, Carr_Kruchkov_2020, Fischer_Klebl_2022]. We treat the middle layer as rigid and relax only the top and bottom layers. More details of the relaxation are presented in SM.
To describe the electronic properties of TTG, we employ a tight-binding model for $p_{z}$ atomic orbitals defined by Eq. (1) and (2). We use the following parameters $a=1.412$ Å, $c=3.36$ Å, $\gamma_{0}=-2.835$ eV, $\gamma_{1}=0.48$ eV. Here, $a$ is the carbon-carbon distance, $c$ is the inter-layer distance of an unrelaxed system, $\gamma_{0}(\gamma_{1})$ is the value of the nearest in-plane (out-of-plane) neighbor hopping. The dimensionless decay constants are $\lambda_{1}=3.15$ in-plane and $\lambda_{2}=7.50$ out-of-plane. A more detailed description of the model can be found in Refs. [TramblydeLaissardiere_Magaud_2010, Rodrigues_Hawrylak_2024].
The considered values of the staggered potential induced by the presence of an hBN substrate were taken to be between 0500-50 meV. The order of these values is consistent with experiments [Hunt_Ashoori_2013, Woods_Novoselov_2014] and ab initio calculations [Giovannetti_Brink_2007, Slawinska_Klusek_2010, Sachs_Lichtenstein_2011, Zollner_Fabian_2019, Zollner_Fabian_2021] for different experimental set-ups, such as aligned graphene/hBN layers, entirely misaligned structure or modulated distance between TBG and the substrate.
@@
[…]
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Funding
A.W.R., D.M. and P.H. were supported by NSERC Discovery Grant No. RGPIN 2019-05714, NSERC Alliance Quantum Grant No. ALLRP/578466-2022, the QSP-078 project of the Quantum Sensors Program at the National Research Council of Canada, University of Ottawa Research Chair in Quantum Theory of Materials, Nanostructures, and Devices. M.B. has been supported by the National Science Centre, Poland, under Grant No. 2021/43/D/ST3/01989.
@@ Acknowledgments
This paper was partly enabled by support provided by the Digital Research Alliance of Canada [AllianceCanada].